Graph the linear equation by finding and plotting its intercepts.
-5x+3y = -15

Answer by rfer(16322)

-5*0+3y=-15
3y=-15
y=-5
(0,-5)
---------------
5x+3*0=-15
5x=-15
x=-3
(-3,0)